Here is a set of data.

86 88 230 371 419 480 489 511 534 706 756 793 889 895


Identify the 5 number summary (min, Q1, median, Q3, max)
____, ____, ____, ____, ____
Answer #1

Min = 86

Q1 = (n+1) / 4 th observation

= (14+ 1) / 4

= 3.75 th observation

= 335.75

Median = (n+1) / 2 th observation

= (14+1) / 2

= 7.5 th observation

= 500

Q3 = 3(n+1) / 4 th observation

= 3( 14 +1) / 4

= 11.25 th observation

= 765.25

Max = 895

5 number summery is

(86 , 335.75 , 500 , 765.25 , 895)